OpenStudy (allymariex3):

Add. (4c2 - 5c + 11) + (3c2 + c - 9) A.7c2 - 4c + 2 B.7c2 + 4c + 20 C.7c2 - 6c + 2 D.7c2 + 6c + 20

OpenStudy (anonymous):

A

OpenStudy (aripotta):

that's correct, all you're doing is combining like terms

OpenStudy (aripotta):

4c2 + 3c2 = 7c2 -5c + c = -4c 11 - 9 = 2

OpenStudy (anonymous):